(D² − 3DD + 2D²)z = e²x+³y +sin(x-2y)

1
Expert's answer
2022-02-04T12:10:34-0500

The given question is;

"\displaystyle
(D^2 − 3DD\prime + 2(D\prime)^2)z = e^{2x+3y}+\sin(x-2y)"

Now put "D=m,\text{and }D\prime=1" then the auxiliary equation is;

"\displaystyle
m^2-3m+2=0\Rightarrow(m-2)(m-1)=0\Rightarrow m=1,2"

Hence, the complementary function is;

"\displaystyle
z_c=f_1(y+x)+f_2(y+2x)".


Next, for the particular integral ("\displaystyle
y_p" );

"\displaystyle
z_p=\frac{1}{D^2 − 3DD\prime + 2(D\prime)^2}(e^{2x+3y}+\sin(x-2y))\\
\quad=\frac{1}{D^2 − 3DD\prime + 2(D\prime)^2}e^{2x+3y}+\frac{1}{D^2 − 3DD\prime + 2(D\prime)^2}\sin(x-2y)\\
\quad=\frac{1}{2^2-3(2)(3)+2(3)^2}e^{2x+3y}+\frac{1}{-(1)^2 − 3[-1\times(-2)] + 2[-(-2)^2]}\sin(x-2y)\\
\quad=\frac{1}{4-18+18}e^{2x+3y}+\left(\frac{1}{-1-3[2]+2[-4]}\sin(x-2y)\right)\\
\quad=\frac{1}{4}e^{2x+3y}+\frac{1}{(-1-6-8)}\sin(x-2y)\\
\quad=\frac{1}{4}e^{2x+3y}+\frac{1}{(-15)}\sin(x-2y)\\
\quad=\frac{1}{4}e^{2x+3y}-\frac{1}{15}\sin(x-2y)"


Thus the general solution is ; "\displaystyle
z=z_c+z_p=f_1(y+x)+f_2(y+2x)+\frac{1}{4}e^{2x+3y}-\frac{1}{15}\sin(x-2y)"